### v1

# Random currents for signed interactions in the one-dimensional heterogeneous Ising model

Abstract.
The random current representation has been very successful in analyzing the ferromagnetic Ising model. In this article, we extend the method to the non-ferromagnetic case. To show its usefulness, we derive a covariance identity for the heterogeneous one-dimensional Ising model, which recovers the correct rate of exponential decay of covariances in the case of random i.i.d. interaction and external field.
Key words and phrases:
Random Ising model, antiferromagnetic, disodered systems, Ising, decay of correlations, random current, transfer matrix
1991 Mathematics Subject Classification:
Primary 82B20; Secondary 60K35
1. Introduction
1.1. The random current representation
The random current representation is an alternative picture of the Ising model, first introduced by Aizenman in [Aiz82] as an extension of ideas of Griffiths, Hurst, and Sherman in [GHS70]. Since then, it has been shown to be remarkably useful in proving a wide range of results, including for the related applications such as the classical XY model [ABF87, ADCS15, DC16, DCT16, Lam23]. To illustrate the motivation for this picture, we consider an Ising system on an arbitrary finite graph $G$ with a collection $S\subseteq 2^{G}$ of subsets of $G$ whose spins interact, and a Hamiltonian
(1.1) $H(\boldsymbol{\sigma})=\sum_{s\in S}j_{s}\prod_{x\in s}\sigma_{x}$
for spin configurations $\boldsymbol{\sigma}\in\{\pm 1\}^{G}$ given the non-negative indexed coefficients $\{j_{s}\geq 0\}_{s\in S}$. It is possible to expand the partition function of the system in the form
(1.2) $Z$ $=\sum_{\boldsymbol{\sigma}\in\{\pm 1\}^{G}}e^{\sum_{s\in S}j_{s}\prod_{x\in s}\sigma_{x}}$
(1.3) $=\sum_{\boldsymbol{\sigma}\in\{\pm 1\}^{G}}\prod_{s\in S}\sum_{n_{s}\in\mathbb{N}_{0}}\frac{j_{s}^{n_{s}}\prod_{x\in s}\sigma_{x}^{n_{s}}}{n_{s}!}$
(1.4) $=\sum_{\mathbf{n}=\{n_{s}\in\mathbb{N}_{0}\}_{s\in S}}\prod_{s\in S}\frac{j_{s}^{n_{s}}}{n_{s}!}\sum_{\boldsymbol{\sigma}\in\{\pm 1\}^{G}}\prod_{x\in s}\sigma_{x}^{n_{s}}.$
Noting the factor of $j_{s}^{n_{s}}/n_{s}!$ that appears in the expansion for each interacting set $s\in S$ given any realization of $\mathbf{n}$, it is possible to interpret $\{n_{s}\}_{s\in S}$ as a set of integer-valued “currents” on the interacting subsets of the lattice distributed as independent Poisson random variables, with respective rates given by $\{j_{s}\}_{s\in S}$. These currents can then be viewed as mediators of correlations between various sites, with the covariance between any two spins being able to be expressed as the probability of an appropriate connection between two sites by these random currents, normalized by the measure of the set of currents that do not have any vertices as sources.
To our knowledge, the random current representation has not been used previously to handle negative interactions, most likely because the convenient interpretation of the interaction coefficients $\{j_{s}\}_{s\in S}$ as rates of Poisson random variables is no longer viable. As a result, the expansion associates to each configuration of currents a signed quantity rather than a probability, making estimates more difficult than in cases with exclusively non-negative interactions. We will be using the model described below in Subsection 1.2 as a test case and show that even in the case of signed interactions the random current representation yields the exact rate of exponential decay in covariance between two spins with distance, in agreement with results found by other methods such as in [CNPV90, CPV12].
1.2. The model
We consider a nearest-neighbor heterogeneous Ising chain, not necessarily ferromagnetic, with an external magnetic field on a one-dimensional finite lattice $\{-N,\ldots,N\}\subset\mathbb{Z}$. The model takes as parameters a set of heterogeneous interaction strengths, denoted by $\mathbf{J}:=\{J_{(x,x+1)}\}_{x=-N}^{N-1}$, as well as a set of field strengths $\mathbf{h}:=\{h_{x}\}_{x=-N}^{N}$. These parameters may take any real values and will be assumed to be distributed according to a probability measure $\nu$. At this point we make no assumptions on $\nu$, and in particular we allow it to be spatially variant or degenerate, unless stated otherwise. Where appropriate, $\mathbb{E}[\cdot]$ will denote the expectation of the enclosed quantity with respect to $\nu$, i.e., over any randomness in the model parameters.
The quenched Hamiltonian for this system is taken to be
(1.5) $H(\boldsymbol{\sigma})=\sum_{x=-N}^{N-1}J_{(x,x+1)}\sigma_{x}\sigma_{x+1}+\sum_{x=-N}^{N}h_{x}\sigma_{x},$
for spin configurations $\boldsymbol{\sigma}:=\{\sigma_{x}\}_{x=-N}^{N}\in\Omega:=\{\pm 1\}^{\{-N,\ldots,N\}}$. The Hamiltonian will also absorb dependence on the inverse temperature in our treatment. The quenched finite-lattice Gibbs measure on the configuration space $\Omega$ is then given by
(1.6) $\mu(\boldsymbol{\sigma})=\frac{e^{H(\boldsymbol{\sigma})}}{Z},$
with the partition function $Z$ defined by
(1.7) $Z=\sum_{\boldsymbol{\sigma^{\prime}}\in\Omega}e^{H(\boldsymbol{\sigma^{\prime}})}.$
[…]
(1.8) $\operatorname{cov}(\sigma_{i},\sigma_{j})=\langle\sigma_{i}\sigma_{j}\rangle-\langle\sigma_{i}\rangle\langle\sigma_{j}\rangle.$
1.3. The transfer matrix method
Historically, the so-called “transfer-matrix method” has proven useful for computing the free energy per site of Ising models in one dimension and on rectangular lattices [CPV12, Ons44]. In one dimension, for a given interval $\{i,\ldots,j\}\subset\mathbb{Z}$ with an appropriate associated Hamiltonian, the partition function for spin configurations on this interval with free boundary conditions can be expressed in the form
(1.9) $Z_{\{i,\ldots,j\}}=\left\|\mathbf{T}_{\{i,\ldots,j\}}\right\|_{\ell^{1}},$
with the Ising transfer matrix product defined
(1.10) $\mathbf{T}_{\{i,\ldots,j\}}:=\begin{bmatrix}e^{h_{j}}&0\\ 0&e^{h_{j}}\end{bmatrix}\prod_{x=i}^{j-1}\begin{bmatrix}e^{J_{(x,x+1)}}&e^{-J_{(x,x+1)}}\\ e^{-J_{(x,x+1)}}&e^{J_{(x,x+1)}}\end{bmatrix}\begin{bmatrix}e^{h_{x}}&0\\ 0&e^{-h_{x}}\end{bmatrix}.$
The exponential growth rate of $\left\|\mathbf{T}_{\{i,\ldots,j\}}\right\|_{\ell^{1}}$ in the thermodynamic limit is thus of interest when computing the free energy per site of the infinite chain and has been solved analytically for homogeneous Ising models on one-dimensional and rectangular lattices [Isi25, Ons44], while for cases with i.i.d. coefficients there exist efficient methods to approximate the free energy per site to arbitrary numerical precision [PS92, PPS95]. Accordingly, results obtained using the transfer matrix method can be used as a benchmark to evaluate the capability of the random current method.
2. Main results
2.1. The truncated chain
As we will be studying the behavior of the quenched covariance between spins located at sites $i,j\in\{-N,\ldots,N\}$, with $i<j$ without loss of generality, it will be useful to define a truncated chain on the vertices $\{i,\ldots,j\}\subset\{-N,\ldots,N\}$, with a modified Hamiltonian
(2.1) $H^{\prime}(\boldsymbol{\sigma}|_{\{i,\ldots,j\}})=\sum_{x=i}^{j-1}J_{(x,x+1)}^{\prime}\sigma_{x}\sigma_{x+1}+\sum_{x=i}^{j}h_{x}^{\prime}\sigma_{x},$
for $\boldsymbol{\sigma}|_{\{i,\ldots,j\}}\in\{\pm 1\}^{\{i,\ldots,j\}}=:\Omega^{\prime}$ where the coefficients $\{J_{(x,x+1)}^{\prime}\}_{x=i}^{j-1}$ and $\{h^{\prime}_{x}\}_{x=i}^{j}$ are chosen such that the measure
(2.2) $\mu^{\prime}(\boldsymbol{\sigma}):=\frac{e^{H^{\prime}(\boldsymbol{\sigma})}}{\sum_{\boldsymbol{\sigma^{\prime}}\in\Omega^{\prime}}e^{H^{\prime}(\boldsymbol{\sigma^{\prime}})}}$
is equal to the restriction of the measure $\mu$ to the $\sigma$-algebra generated by $\{\sigma_{i},\ldots,\sigma_{j}\}$. One can always achieve this in a way such that $J^{\prime}_{(x,x+1)}=J_{(x,x+1)}$ for all $i\leq x<j$ and that $h^{\prime}_{x}=h_{x}$ for all $i<x<j$, with the only coefficients that differ from the original model being $h^{\prime}_{i}$ and $h^{\prime}_{j}$; at those sites, moreover, we have the bounds $|h^{\prime}_{i}-h_{i}|\leq|J_{(i-1,i)}|$ and $|h^{\prime}_{j}-h_{j}|\leq|J_{(j,j+1)}|$. We refer the interested reader to Section A in the appendix for the precise calculations.
To distinguish this truncated chain from the original one, we will refer to the chain on $\{i,\ldots,j\}$ as the truncated chain and the chain on $\{-N,\ldots,N\}$ as the long chain.
2.2. Statement of main results
As our main result, we illustrate that random currents can be used to derive the following matrix identity for the covariance, regardless of the signs of the interaction coefficients.
Theorem 2.1.
For the one-dimensional truncated chain described above,
(2.3) $|\operatorname{cov}(\sigma_{i},\sigma_{j})|=\frac{1}{(e_{1}^{\intercal}Me_{1})^{2}},$
where $e_{1}$ is the first standard basis vector in $\mathbb{R}^{2}$, and
(2.4) $\begin{split}M&:=\begin{bmatrix}\cosh(h^{\prime}_{j})&\sinh(h^{\prime}_{j})\\ \sinh(h^{\prime}_{j})&\cosh(h^{\prime}_{j})\end{bmatrix}\\ &\qquad\cdot\prod_{x=i+1}^{j-1}\left\{\begin{bmatrix}\sqrt{\coth|J_{(x,x+1)}|}&0\\ 0&\operatorname{sign}(J_{(x,x+1)})\sqrt{\tanh|J_{(x,x+1)}|}\end{bmatrix}\begin{bmatrix}\cosh(h_{x})&\sinh(h_{x})\\ \sinh(h_{x})&\cosh(h_{x})\end{bmatrix}\right\}\\ &\qquad\cdot\begin{bmatrix}\sqrt{\coth|J_{(i,i+1)}|}&0\\ 0&\operatorname{sign}(J_{(i,i+1)})\sqrt{\tanh|J_{(i,i+1)}|}\end{bmatrix}\begin{bmatrix}\cosh(h^{\prime}_{i})&\sinh(h^{\prime}_{i})\\ \sinh(h^{\prime}_{i})&\cosh(h^{\prime}_{i})\end{bmatrix}.\end{split}$
Remark 2.2.
Under a change of basis, using $\frac{1}{\sqrt{2}}(e_{1}+e_{2})$ and $\frac{1}{\sqrt{2}}(e_{1}-e_{2})$ as the new basis vectors, the matrix in (2.4) is equivalent to a normalized Ising transfer matrix product
(2.5) $\begin{split}\tilde{\mathbf{T}}_{\{i\cdots j\}}:=\begin{bmatrix}e^{h_{j}}&0\\ 0&e^{-h_{j}}\end{bmatrix}\prod_{x=i}^{j-1}\frac{1}{\sqrt{2\sinh(2|J_{x,x+1}|)}}\begin{bmatrix}e^{J_{x,x+1}}&e^{-J_{x,x+1}}\\ e^{-J_{x,x+1}}&e^{J_{x,x+1}}\end{bmatrix}\begin{bmatrix}e^{h_{x}}&0\\ 0&e^{-h_{x}}\end{bmatrix}\\ \cdot\frac{1}{\sqrt{2\sinh(2|J_{(i,i+1)}|)}}\begin{bmatrix}e^{J_{(i-i+1)}}&e^{-J_{(i,i+1)}}\\ e^{-J_{(i,i+1)}}&e^{J_{(i,i+1)}}\end{bmatrix}\begin{bmatrix}e^{h_{i}}&0\\ 0&e^{-h_{i}}\end{bmatrix},\end{split}$
a rescaling of the original transfer matrix product analogous to the matrix product for the long chain in (1.10) that has determinant ±$\pm 1$.
It should be noted that Theorem 2.4 can be deduced using other methods, and we refer the interested reader to Section B in the appendix for a proof of this identity using standard techniques of statistical mechanics. Although we are not aware of any sources in which it has been stated, we believe that it is not new due to its connection to Ising’s transfer matrix method. Nevertheless, it does demonstrate that such an identity expressing covariance can be obtained using random currents on a model with signed coefficients. To illustrate the utility of Theorem 2.4 in understanding the behavior of covariances, we also prove a few supplemental results quantifying the decay in covariances. First, we prove a lower bound on the absolute value of the covariance in all cases.
Corollary 2.3.
For the truncated chain $\{i,\ldots,j\}$,
(2.6) $|\operatorname{cov}(\sigma_{i},\sigma_{j})|\geq\left(\prod_{x=i}^{j-1}\tanh|J_{(x,x+1)}|\right)e^{-2(|h^{\prime}_{i}|+\sum_{x=i+1}^{j-1}|h_{x}|+|h^{\prime}_{j}|)}.$
It follows that in terms of the original coefficients, we have the bound
(2.7) $|\operatorname{cov}(\sigma_{i},\sigma_{j})|\geq\left(\prod_{x=i}^{j-1}\tanh|J_{(x,x+1)}|\right)e^{-2(\sum_{x=i}^{j}|h_{x}|+|J_{(i-1,i)}|+|J_{(j,j+1)}|)}.$
In the case where $\nu$ on the space of non-negative real values for all $\{J_{(x,x+1)}\}_{x=-N}^{N-1}$ and $\{h_{x}\}_{x=-N}^{N}$, we additionally prove the following upper bound on covariance.
Corollary 2.4.
If $(\mathbf{J},\mathbf{h})$ are all non-negative, then for the truncated chain, we have
(2.8) $|\operatorname{cov}(\sigma_{i},\sigma_{j})|\leq\left(\prod_{x=i}^{j-1}\frac{\tanh(J_{(x,x+1)})}{(1+e^{-2J_{(x,x+1)}})^{2}}\right)\frac{1}{\cosh^{2}\left({h^{\prime}_{i}+\sum_{x=i+1}^{j-1}h_{x}+h^{\prime}_{j}}\right)}.$
In terms of the original coefficients, this implies the weaker bound
(2.9) $|\operatorname{cov}(\sigma_{i},\sigma_{j})|\leq\left(\prod_{x=i}^{j-1}\frac{\tanh(J_{(x,x+1)})}{(1+e^{-2J_{(x,x+1)}})^{2}}\right)\frac{1}{\cosh^{2}\left({\sum_{x=i}^{j}h_{x}}\right)}.$
In the limit where $\operatorname{dist}(i,j)$ is large and the coefficients $\mathbf{J}$ and $\mathbf{h}$ are chosen independently and each identically distributed, it becomes insightful to study the product of matrices that appears in (2.4) as a dynamical system [DH81, DH83, CNPV90, PS92, PPS95, GG22]. For convenience, in this case we can consider $\nu$ to be a restriction of a measure that chooses i.i.d. coefficients on all points and nearest-neighbor edges of $\mathbb{Z}$, which we can also denote $\nu$ by abuse of notation. Because the exact values of quenched covariances will depend slightly on where we restrict the lattice, we will use the notation $\operatorname{cov}_{\{-N,\ldots,N\}}(\cdot)$ where necessary to specify the lattice size.
As long as $\nu$ is sufficiently well-behaved that $\mathbb{E}[|\log\left\|M(J,h)\right\||]<\infty$, where
(2.10) $M(J,h):=\begin{bmatrix}\sqrt{\coth|J|}&0\\ 0&\operatorname{sign}(J)\sqrt{\tanh|J|}\end{bmatrix}\begin{bmatrix}\cosh(h)&\sinh(h)\\ \sinh(h)&\cosh(h)\end{bmatrix},$
with $J$ and $h$ adhering to the marginal distributions of $J_{(0,1)}$ and $h_{0}$, respectively, the limit
(2.11) $\gamma:=\lim_{N\to\infty}\frac{\mathbb{E}[\log\left\|M_{N}\cdots M_{1}\right\|]}{N}$
exists for the product of matrices $\{M_{n}\}_{n\in\mathbb{N}}$ sampled from this distribution as a consequence of Fekete’s lemma [Fek23]. The limit in (2.11) is called the upper Lyapunov exponent of the system. Moreover, the ergodicity of $\nu$ implies that this statement can be strengthened so that
(2.12) $\gamma:=\lim_{N\to\infty}\frac{\log\left\|M_{N}\cdots M_{1}\right\|}{N}$
for $\nu$-almost all realizations of $\{M_{n}\}_{n\in\mathbb{N}}$ instead of only in expectation [Fur63, Ose65, Kin73, Lig85]. As a result, the upper Lyapunov exponent of the matrix product characterizes the growth of the expression in (2.4) for almost all cases.
In a loose sense, we can then conclude that covariances should decay asymptotically with a rate given by $|\operatorname{cov}(\sigma_{i},\sigma_{j})|\sim e^{-2\gamma\operatorname{dist}(i,j)}$ as $\operatorname{dist}(i,j)$ grows large. There are, however, two complications that require consideration when we state this property more precisely. The first is that as physical lattices are finite, taking $\operatorname{dist}(i,j)$ to infinity is possible only if we consider a family of lattices of increasing size. Secondly, this convergence will not necessarily be uniform in $(\mathbf{J},\mathbf{h})$ for any full-measure subspace, as a given realization of $(\mathbf{J},\mathbf{h})$ may have outliers on certain edges or sites that causes an arbitrarily large deviation from “typical” behavior for finite lattices. To account for these complications, we can take the limit in the following manner: if a rule $\nu$ for choosing the interaction parameters is fixed, we begin with some lattice $\{-N,\ldots,N\}$ and sites $i$ and $j$. If we add new sites to the ends of the lattice while leaving the parameters on the original sublattice unchanged, and we move $i$ and $j$ further apart within the expanding lattice so that their distance approaches infinity, then we should observe convergence to the exponential decay rate of $-2\gamma$. We state this result in the following corollary.
Corollary 2.5.
If $\mathbb{E}|J_{(0,1)}|<\infty$, $\mathbb{E}[\log(\coth|J_{(0,1)}|)]<\infty$ and $\mathbb{E}|h_{0}|<\infty$, then there exists a set $E$ of coefficients $\{J_{(x,x+1)}\}_{x\in\mathbb{Z}}$ and $\{h_{x}\}_{x\in\mathbb{Z}}$ with $\nu(E)=1$, such that if we select a fixed ($(\{J_{(x,x+1)}\}_{x\in\mathbb{Z}},\{h_{x}\}_{x\in\mathbb{Z}})\in E$ and consider the family of $\{N;i,j\}$ denoting finite sublattices $\{-N,\ldots,N\}$ for $N\in\mathbb{N}$ and points $i<j$ contained in $\{-N,\ldots,N\}$, with coefficients taken by restricting ($(\{J_{(x,x+1)}\}_{x\in\mathbb{Z}},\{h_{x}\}_{x\in\mathbb{Z}})$ to the sublattice, and a partial order such that $\{N;i,j\}\leq\{N^{\prime};i^{\prime},j^{\prime}\}$ if and only if $i^{\prime}\leq i<j\leq j^{\prime}$ then the net
(2.13) $\alpha_{N;i,j}:=\frac{\log|\operatorname{cov}_{\{-N,\ldots,N\}}(\sigma_{i},\sigma_{j})|}{\operatorname{dist}(i,j)}$
has a unique limit point at $-2\gamma$, where $\gamma$ is the upper Lyapunov exponent of the product of matrices defined in (2.11).
Remark 2.6.
The regularity condition $\mathbb{E}[\log(\coth|J_{(0,1)}|)]<\infty$ is necessary to guarantee that nearest-neighbor interactions are strong enough to exhibit exponential decay. Without this condition, it is possible to see superexponential decay, or in extreme cases, for the covariance to drop to zero at finite distance with nonzero probability.
The upper Lyapunov exponent in Corollary 2.5 can be related to the upper Lyapunov exponent of the matrix in (1.10) via Remark 2.2 and elementary calculation. It may not be possible to express the value of this exponent in closed form. However, its behavior has been studied in detail and is possible to compute numerically to arbitrary precision with efficient algorithms [CNPV90, CPV12, DH81, DH83, PPS95, GG22].
3. Random current expansion for signed interactions
In this section, we will consider an arbitrary finite graph $G$, with $S\subseteq 2^{G}$ being the collection of subsets whose spins interact. We write the Hamiltonian in the abbreviated form
(3.1) $H(\boldsymbol{\sigma})=\sum_{s\in S}j_{s}\sigma_{s},$
[…]
(3.2) $Z\left<\sigma_{A}\right>$ $=\sum_{\boldsymbol{\sigma}\in\Omega}\sigma_{A}\exp\left(\sum_{s\in S}j_{s}\sigma_{s}\right)$
(3.3) $=\sum_{\boldsymbol{\sigma}\in\Omega}\sigma_{A}\prod_{s\in S}\sum_{n_{s}\in\mathbb{N}_{0}}\frac{j_{s}^{n_{s}}}{n_{s}!}\sigma_{S}^{n_{s}}$
(3.4) $=\sum_{\mathbf{n}=\{n_{s}\}_{s\in S}}\prod_{s\in S}\frac{j_{s}^{n_{s}}}{n_{s}!}\sum_{\boldsymbol{\sigma}\in\Omega}\sigma_{A}\prod_{s\in S}\sigma_{s}^{n_{s}}$
The collection of indices $\mathbf{n}$ that arises in line (3.4) will be referred to as a current. Moreover, the fact that each current that appears in line (3.4) is weighted by a factor of $\prod_{s\in S}\frac{j_{s}^{n_{s}}}{n_{s}!}$ suggests that we define a signed measure $\mathfrak{P}$ on $\mathcal{C}$, which can be decomposed into the product of individual signed measures each defined on the space of currents on a single element of $S$ (i.e., $\mathbb{N}_{0}$), with
(3.5) $\mathfrak{P}(\mathbf{a})=\prod_{s\in S}\mathfrak{P}_{s}(a_{s}):=\prod_{s\in S}\frac{\exp(-j_{s})j_{s}^{a_{s}}}{a_{s}!}.$
The exponential factor in the definition is included for normalization purposes, so that each $\mathfrak{P}_{s}$, $s\in S$, has the property that $\mathfrak{P}_{s}(\mathbb{N}_{0})=1$. Note that because $j_{s}$ can be negative for a given $s\in S$, this normalization does not make $\mathfrak{P}$ a probability, and hence the name “random current” is a slight misnomer in this general case. Nevertheless, we can define an appropriate extension of the term as follows.
Definition 3.1.
A random current on $G$ is a function $\mathbf{n}$, taking values in $\mathcal{C}:=\mathbb{N}_{0}^{S}$ equipped with the signed measure $\mathfrak{P}$ defined in (3.5).
Remark 3.2.
If $\mathbf{j}$ has exclusively non-negative entries, then $\mathfrak{P}$ is a probability measure, and the entries of a random current with rates $\mathbf{j}$ are independent Poisson random variables with $n_{s}$ having rate $j_{s}$ for every $s\in S$.
Notation 3.3.
We will use the symbols $\mathfrak{P}$ and $\mathfrak{E}$ to denote, respectively, the signed measure of a random current with rates $\mathbf{j}$ and the integral with respect to the signed measure. For the product measure for two currents $\mathbf{n}_{1},\mathbf{n}_{2}\in\mathcal{C}$, we will use the notation $\mathfrak{P}\times\mathfrak{P}$.
Continuing from line (3.4), we then have that
(3.6) $Z\left<\sigma_{A}\right>$ $=\sum_{\mathbf{n}=\{n_{s}\}_{s\in S}}\exp\left(\sum_{s\in S}j_{s}\right)\mathfrak{P}(\mathbf{n})\sum_{\boldsymbol{\sigma}\in\{\pm 1\}^{G}}\prod_{x\in G}\sigma_{x}^{\sum_{s\in S:x\in s}n_{s}+\mathbbm{1}_{\{x\in A\}}}$
(3.7) $=\sum_{\mathbf{n}=\{n_{s}\}_{s\in S}}\exp\left(\sum_{s\in S}j_{s}\right)\mathfrak{P}(\mathbf{n})\prod_{x\in G}\sum_{\sigma_{x}\in\{\pm 1\}}\sigma_{x}^{\sum_{s\in S:x\in s}n_{s}+\mathbbm{1}_{\{x\in A\}}}.$
In the final expression, we observe that $\sum_{\sigma_{i}=\{\pm 1\}}\sigma_{i}^{m_{i}}$ for $m_{i}\in\mathbb{N}_{0}$ is $2$ if $m_{i}$ is even and zero if $m_{i}$ is odd. Therefore,
(3.8) $\prod_{i\in G}\sum_{\sigma_{i}\in\{\pm 1\}}\sigma_{i}^{m_{i}}=\begin{cases}2^{|G|},&\text{$m_{i}$ even for every $i\in G$},\\ 0,&\text{otherwise}.\end{cases}$
For our purposes, we are interested in setting
(3.9) $m_{x}:=\sum_{s\in S:x\in s}n_{s}+\mathbbm{1}_{\{x\in A\}}.$
Thus, it is convenient to make the following definition.
Definition 3.4.
For a current $\mathbf{n}\in\mathcal{C}$, the boundary $\partial\mathbf{n}$ is the set of sites in $G$ given by
(3.10) $\partial\mathbf{n}:=\left\{x\in G:\sum_{s\in S:x\in s}n_{s}\;\mathrm{is\;odd}\right\}.$
With this definition, combining equations (3.7) and (3.8) enable us to compactly write
(3.11) $Z\left<\sigma_{A}\right>=2^{|G|}\exp\left(\sum_{s\in S}j_{s}\right)\mathfrak{P}(\{\mathbf{n}\in\mathcal{C}:\partial\mathbf{n}=A\}).$
It follows that
[…]
(3.13) $=\frac{\mathfrak{P}(\{\mathbf{n}\in\mathcal{C}:\partial\mathbf{n}=A\})}{\mathfrak{P}(\{\mathbf{n}\in\mathcal{C}:\partial\mathbf{n}=\varnothing\})}.$
As a result, we can express the single-site expectation at a vertex $x\in G$ as
(3.14) $\sigma_{x}\right>=\frac{\mathfrak{P}(\{\mathbf{n}\in\mathcal{C}:\partial\mathbf{n}=\{x\}\})}{\mathfrak{P}(\{\mathbf{n}\in\mathcal{C}:\partial\mathbf{n}=\varnothing\})}$
and the mixed second moment in the spins at vertices $x,y\in G$ as
(3.15) $\sigma_{x}\sigma_{y}\right>=\frac{\mathfrak{P}(\{\mathbf{n}\in\mathcal{C}:\partial\mathbf{n}=\{x,y\}\})}{\mathfrak{P}(\{\mathbf{n}\in\mathcal{C}:\partial\mathbf{n}=\varnothing\})}.$
To express the covariance between two spins with respect to the finite-volume Gibbs measure in the language of random currents, we invoke a generalization of the so-called “switching lemma” to allow negative interaction coefficients. The original form of the switching lemma was used in [GHS70, Lemma 1] to prove the GHS inequality, and it was subsequently reformulated by [Aiz82, Lemma 3.2] to apply to random currents.
Lemma 3.5 (Switching lemma [GHS70, Aiz82]).
Let $G$ be a graph, and let $\mathfrak{P}$ be the signed measure for random currents on $G$ with rates $\mathbf{j}=\{j_{s}\}_{s\in S}$, where $S\subset G$ is the set of graph elements that are allowed to accommodate nonzero currents. If $A$ and $B$ are two subsets of $G$, then
(3.16) $\begin{split}\mathfrak{P}&(\{\partial\mathbf{n}_{1}=A\})\cdot\mathfrak{P}(\{\partial\mathbf{n}_{2}=B\})\\ &=\mathfrak{P}\times\mathfrak{P}(\{\partial\mathbf{n}_{1}=\varnothing,\,\partial\mathbf{n}_{2}=A\triangle B,\,\partial\mathbf{k}=A\;\mathrm{for\;some}\;\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}\}).\end{split}$
The proof of the lemma involves a reduction from the statement about random currents to the purely combinatorial statement first found in [GHS70], which can then be verified by an elementary counting argument. We review the proof here, as its mechanism is important to justify the application of the lemma to random currents with negative interactions.
Proof.
Recall that for a single current $\mathbf{n}$,
(3.17) $\mathfrak{P}(\mathbf{n})=\prod_{s\in S}\frac{e^{-j_{s}}j_{s}^{n(s)}}{n(s)!}.$
This signed measure exhibits the multiplication law
(3.18) $\mathfrak{P}(\mathbf{n}_{1})\mathfrak{P}(\mathbf{n}_{2})$ $=\prod_{s\in S}\frac{\left(e^{-j_{s}}\right)^{2}j_{s}^{n_{1}(s)+n_{2}(s)}}{n_{1}(s)!n_{2}(s)!}$
(3.19) $=\prod_{s\in S}\frac{\left(e^{-j_{s}}\right)^{2}j_{s}^{n_{1}(s)+n_{2}(s)}}{(n_{1}(s)+n_{2}(s))!}\binom{n_{1}(s)+n_{2}(s)}{n_{1}(s)}$
(3.20) $=e^{-\sum_{s\in S}j_{s}}\mathfrak{P}(\mathbf{n}_{1}+\mathbf{n}_{2})\binom{\mathbf{n}_{1}+\mathbf{n}_{2}}{\mathbf{n}_{1}},$
where as a shorthand, we use the convention that
(3.21) $\binom{\mathbf{n}_{1}+\mathbf{n}_{2}}{\mathbf{n}_{1}}:=\prod_{s\in S}\binom{n_{1}(s)+n_{2}(s)}{n_{1}(s)}.$
As a result, the expression on the left-hand side of (3.16) can be rewritten in the form
(3.22) $\mathfrak{P}(\{\partial\mathbf{n}_{1}=A\})\cdot\mathfrak{P}(\{\partial\mathbf{n}_{2}=B\})$ $=\sum_{\partial\mathbf{n}_{1}=A}\sum_{\partial\mathbf{n}_{2}=B}\mathfrak{P}(\mathbf{n}_{1})\mathfrak{P}(\mathbf{n}_{2})$
(3.23) $=e^{-\sum_{s\in S}j_{s}}\sum_{\partial\mathbf{n}_{1}=A}\sum_{\partial\mathbf{n}_{2}=B}\mathfrak{P}(\mathbf{n}_{1}+\mathbf{n}_{2})\binom{\mathbf{n}_{1}+\mathbf{n}_{2}}{\mathbf{n}_{1}}.$
Making the variable change $\mathbf{m}:=\mathbf{n}_{1}+\mathbf{n}_{2}$ and $\mathbf{n}:=\mathbf{n}_{1}$, the range of summation becomes $\{\partial\mathbf{m}=A\triangle B,\partial\mathbf{n}=A,\mathbf{n}\leq\mathbf{m}\}$. Thus,
(3.24) $\mathfrak{P}(\{\partial\mathbf{n}_{1}=A\})\cdot\mathfrak{P}(\{\partial\mathbf{n}_{2}=B\})=e^{-\sum_{s\in S}j_{s}}\sum_{\partial\mathbf{m}=A\triangle B}\mathfrak{P}(\mathbf{m})\sum_{\partial\mathbf{n}=A,\,\mathbf{n}\leq\mathbf{m}}\binom{\mathbf{m}}{\mathbf{n}}.$
On the right-hand side, the expression can be rewritten as
[…]
(3.25) $=\sum_{\partial\mathbf{n}_{1}=\varnothing}\sum_{\partial\mathbf{n}_{2}=A\triangle B}\mathfrak{P}(\mathbf{n}_{1})\mathfrak{P}(\mathbf{n}_{2})\mathbbm{1}_{\{\partial\mathbf{k}=A\,\text{for some}\,\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}\}}$
(3.26) $=e^{-\sum_{s\in S}j_{s}}\sum_{\partial\mathbf{m}=A\triangle B}\mathfrak{P}(\mathbf{m})\sum_{\partial\mathbf{n}=\varnothing,\mathbf{n}\leq\mathbf{m}}\binom{\mathbf{m}}{\mathbf{n}}\mathbbm{1}_{\{\partial\mathbf{k}=A\,\text{for some}\,\mathbf{k}\leq\mathbf{m}\}},$
where we made the same transformation to the new currents $\mathbf{m}$ and $\mathbf{n}$.
Comparison of the expressions (3.24) and (3.26) now reveals that it suffices to show that, for any current $\mathbf{m}$ such that $\partial\mathbf{m}=A\triangle B$,
(3.27) $\sum_{\partial\mathbf{n}=A,\,\mathbf{n}\leq\mathbf{m}}\binom{\mathbf{m}}{\mathbf{n}}=\begin{cases}\sum_{\partial\mathbf{n}=\varnothing,\mathbf{n}\leq\mathbf{m}}\binom{\mathbf{m}}{\mathbf{n}},&\partial\mathbf{k}=A\,\text{for some}\,\mathbf{k}\leq\mathbf{m},\\ 0,&\text{otherwise}.\end{cases}$
To prove this statement, we will initialize an empty family of sets $\mathcal{M}$. For each $s\in S$, we will take $m_{s}$ copies of the set $s$ and add them to $\mathcal{M}$. For convenience, we will let the operator $\Delta\mathcal{N}$, acting on subfamilies $\mathcal{N}\subset\mathcal{M}$, denote the symmetric difference of all sets in $\mathcal{N}$. Then, the left-hand side of (3.27) counts the number of subfamilies $\mathcal{N}\subset\mathcal{M}$ such that $\Delta\mathcal{N}=A$. On the right-hand side, in the first case we have by assumption a fixed subfamily $\mathcal{K}\subset\mathcal{M}$ such that $\Delta\mathcal{K}=A$. Under this condition, it can be seen that the map $\mathcal{N}\mapsto\mathcal{N}\triangle\mathcal{K}$ is a bijection between the subfamilies $\{\mathcal{N}\subset\mathcal{M}:\Delta\mathcal{N}=A\}$ and $\{N\subset M:\Delta\mathcal{N}=\varnothing\}$. As a result, the two collections of subfamilies have equal cardinality, so equality holds. The second case becomes tautological as the left-hand side of (3.27) becomes an empty summation by assumption. ∎
Crucially for our purposes, the reduction from (3.16) to (3.27) required no conditions on the signs of $\mathbf{j}$, and (3.27) no longer exhibits any dependence on $\mathbf{j}$ at all. A direct corollary of this lemma is the following expression for covariance.
Corollary 3.6.
Suppose that $x$ and $y$ are vertices in the lattice of an Ising model and $\mathfrak{P}$ again denotes the signed measure for random currents on the lattice. Then,
(3.28) $\operatorname{cov}(\sigma_{i},\sigma_{j})=\frac{\mathfrak{P}\times\mathfrak{P}(\{\partial\mathbf{n}_{1}=\varnothing;\partial\mathbf{n}_{2}=\{i,j\};\text{$\partial\mathbf{k}\neq\{i\}$ for any $\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}$}\})}{\mathfrak{P}(\{\partial\mathbf{n}=\varnothing\})^{2}}.$
Proof.
This follows immediately from Equations (3.14) and (3.15) with the application of Lemma 3.16. ∎
4. Application to the truncated chain
While the events in the right-hand expression in (3.28) can be rather complicated for general lattices, the uniqueness of paths between points in a one-dimensional lattice makes it possible to simplify the equation into an explicit formula for the truncated chain, eventually yielding Theorem 2.4.
Notation 4.1.
In this section, we will let $V$ refer to the set of points $\{i,\ldots,j\}$ in the truncated chain and $E$ refer to the set of nearest-neighbor edges $\{(i,i+1),\ldots,(j-1,j)\}$. Moreover, $\mathcal{C}^{V}$ will denote the space of currents supported on $V$, and $\mathcal{C}^{E}$ will denote the space of currents supported on $E$. Any random current $\mathbf{n}\in\mathbf{N}_{0}^{V\cup E}$ on the truncated chain can be decomposed uniquely in the form $\mathbf{n}=\mathbf{n}^{V}\oplus\mathbf{n}^{E}$, where $\mathbf{n}^{V}\in\mathcal{C}^{V}$ and $\mathbf{n}^{E}\in\mathcal{C}^{E}$.
We will utilize this path-uniqueness by first conditioning the events in (3.28) on $\mathbf{n}_{1}^{V}$, $\mathbf{n}_{2}^{V}$, and $\mathbf{n}^{V}$. Because the dependence of $\mathbf{n}_{1}^{E}$, $\mathbf{n}_{2}^{E}$, and $\mathbf{n}^{E}$ on these variables will be global, we introduce the following terminology for simplicity:
Definition 4.2.
If $\mathbf{n}^{V}$ is a current on $V$, then we will say that an edge $e=(x,x+1)\in E$ “splits $\mathbf{n}^{V}$ evenly” if $\sum_{y\leq x}\mathbf{n}^{V}(y)$ and $\sum_{y\geq x}\mathbf{n}^{V}(y)$ are even and that $e$ “splits $\mathbf{n}^{V}$ oddly” if both summations are odd. If $\boldsymbol{\epsilon}=\{\epsilon_{e}\}_{e\in E}\in\{\mathtt{odd},\mathtt{even}\}^{E}$, then we will say that $\mathbf{n}^{V}$ is split according to $\boldsymbol{\epsilon}$ if individual edges $e\in E$ split $\mathbf{n}^{V}$ evenly whenever $\epsilon_{e}=\mathtt{even}$ and oddly whenever $\epsilon_{e}=\mathtt{odd}$.
We will show that evaluating the conditional measures yields the following expression for covariances.
Lemma 4.3.
If the underlying lattice in (3.28) is the truncated chain, then we have the identity
(4.1) $|\operatorname{cov}(\sigma_{i},\sigma_{j})|=\left(\frac{e^{-\sum_{i\leq x\leq j}h_{x}}}{\mathfrak{E}\left[\prod_{e\in E}f_{e}(\mathbf{n}^{V})\right]}\right)^{2},$
where
(4.2) $f_{e}(\mathbf{n}^{V}):=\sqrt{\coth|J_{e}|}\mathbbm{1}_{\text{$e$ \rm{splits} $\mathbf{n}_{V}$ \rm{evenly}}}+\operatorname{sign}(J_{e})\sqrt{\tanh|J_{e}|}\mathbbm{1}_{\text{$e$ \rm{splits} $\mathbf{n}_{V}$ \rm{oddly}}},$
and the integral in the denominator of (4.1) is taken over all $\mathbf{n}^{V}\in\mathcal{C}^{V}$.
As the first step in proving this lemma, we first observe that the numerator becomes particularly simple due to path-uniqueness, with conditioning on $\mathbf{n}_{1}^{V}$ and $\mathbf{n}_{2}^{V}$ becoming trivial.
Lemma 4.4.
If the underlying lattice in (3.28) is one-dimensional with vertices $i$ and $j$ being opposite endpoints, then the event in the numerator for $\mathbf{n}_{1},\mathbf{n}_{2}\in\mathcal{C}$ is equivalent to the following three conditions all being met:
  1. (1)
    $\mathbf{n}_{1}^{E}(e)$ is even for all $e\in E$;
  2. (2)
    $\mathbf{n}_{2}^{E}(e)$ is odd for all $e\in E$; and
  3. (3)
    $\mathbf{n}_{1}^{V}$ and $\mathbf{n}_{2}^{V}$ are both identically zero.
As a result,
(4.3) $\begin{split}|\mathfrak{P}\times\mathfrak{P}(\{&\partial\mathbf{n}_{1}=\varnothing;\partial\mathbf{n}_{2}=\{i,j\};\text{$\partial\mathbf{k}\neq\{i\}$\rm{for any} $\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}$}\})|\\ &=\left(e^{-\sum_{x\in V}h_{x}}\prod_{e\in E}\sqrt{\mathfrak{P}(\mathbf{n}^{E}(e)\,\mathrm{even})\mathfrak{P}(\mathbf{n}^{E}(e)\,\mathrm{odd})}\right)^{2}.\end{split}$
Proof.
First, suppose that $(\mathbf{n}_{1},\mathbf{n}_{2})\in\mathcal{C}^{2}$ belongs to the event in the numerator of (3.28). The condition that $\partial\mathbf{k}\neq\{i\}$ for any $\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}$ implies that $\mathbf{n}_{1}^{V}(i)=\mathbf{n}_{2}^{V}(i)=0$, and $\mathbf{n}_{1}^{V}(x)=\mathbf{n}_{2}^{V}(x)=0$ for any $x$ that is connected to $i$ via $\mathbf{n}_{1}+\mathbf{n}_{2}$.
Using these observations and considering the requirements that $i\notin\partial\mathbf{n}_{1}$ and that $i\in\partial\mathbf{n}_{2}$, we deduce that $\mathbf{n}_{1}^{E}(i,i+1)$ is even and that $\mathbf{n}_{2}^{E}(i,i+1)$ is odd. Thus, $i+1$ is connected to $i$ at least by $\mathbf{n}_{2}$, and so $\mathbf{n}_{1}^{V}(i+1)=\mathbf{n}_{2}^{V}(i+1)=0$. Proceeding inductively, if vertex $x>i$ is connected to $i$ with $\mathbf{n}_{1}^{E}(x-1,x)$ being even and $\mathbf{n}_{2}^{E}(x-1,x)$ being odd, then $\mathbf{n}_{1}^{V}(x)=\mathbf{n}_{2}^{V}(x)=0$, and if $x<j$ then the fact that $x\notin\partial\mathbf{n}_{1}\cup\partial\mathbf{n}_{2}$ implies that $\mathbf{n}_{1}^{E}(x,x+1)$ is even while $\mathbf{n}_{2}^{E}(x,x+1)$ is odd and connects $x+1$ to all vertices to its left. This recursion continues until it reaches $j$, verifying that the conditions listed in Lemma 4.3 are all met.
Conversely, suppose that all conditions in the lemma are satisfied. By examining the outgoing currents at any vertex, it is easy to verify that $\partial\mathbf{n}_{1}=\varnothing$ and that $\partial\mathbf{n}_{2}=\{i,j\}$. Moreover, given that $\mathbf{n}_{1}^{V}$ and $\mathbf{n}_{2}^{V}$ are identically zero, we have that $\mathbf{k}=\mathbf{k}^{E}$ for any $\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}$, and incrementing $\mathbf{k}^{E}$ along any edge changes #$\#(\partial\mathbf{k})$ by an even number. Because the boundary of the zero current is empty and every $\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}$ can be obtained by starting with the zero current and incrementing its value along edges a finite number of times, we can conclude that $\partial\mathbf{k}=\{i\}$ is impossible for any $\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}$, so that $(\mathbf{n}_{1},\mathbf{n}_{2})$ belongs to the event in the numerator of (3.28).
Thus,
(4.4) $\begin{split}|\mathfrak{P}\times\mathfrak{P}(\{&\partial\mathbf{n}_{1}=\varnothing;\partial\mathbf{n}_{2}=\{i,j\};\text{$\partial\mathbf{k}\neq\{i\}$ for any $\mathbf{k}\leq\mathbf{n}_{1}+\mathbf{n}_{2}$}\})|\\ &=|\mathfrak{P}(\mathbf{n}_{1}^{E}\,\mathrm{even})\mathfrak{P}(\mathbf{n}_{2}^{E}\,\mathrm{odd})\mathfrak{P}(\mathbf{n}_{1}^{V}\equiv 0)\mathfrak{P}(\mathbf{n}_{2}^{V}\equiv 0)|\end{split}$
(4.5) $=\prod_{e\in E}|\mathfrak{P}(\mathbf{n}^{E}(e)\,\text{even})|\cdot\prod_{e\in E}|\mathfrak{P}(\mathbf{n}^{E}(e)\,\text{odd})|\cdot\left(e^{-\sum_{x\in V}h_{x}}\right)^{2}$
(4.6) $=\left(e^{-\sum_{x\in V}h_{x}}\prod_{e\in E}\sqrt{|\mathfrak{P}(\mathbf{n}^{E}(e)\,\text{even})\mathfrak{P}(\mathbf{n}^{E}(e)\,\text{odd})|}\right)^{2},$
verifying (4.3). ∎
The denominator of (3.28) does not similarly simplify, but it can be reduced in terms of an event involving only one random current, since
(4.7) $\mathfrak{P}((\mathbf{n}_{1},\mathbf{n}_{2}):\partial\mathbf{n}_{1}=\varnothing;\partial\mathbf{n}_{2}=\varnothing)=\mathfrak{P}(\mathbf{n}:\partial\mathbf{n}=\varnothing)^{2}.$
With the decomposition $\mathbf{n}=\mathbf{n}^{V}\oplus\mathbf{n}^{E}$, the condition $\partial\mathbf{n}=\varnothing$ is equivalent to $\partial\mathbf{n}^{E}=\partial\mathbf{n}^{V}$. As a result, conditioning on $\mathbf{n}^{V}$ yields that
(4.8) $\mathfrak{P}(\mathbf{n}:\partial\mathbf{n}=\varnothing)$ $=\mathfrak{P}(\mathbf{n}:\partial\mathbf{n}^{E}=\partial\mathbf{n}^{V})$
(4.9) $=\mathfrak{E}(\mathfrak{P}(\partial\mathbf{n}^{E}=\partial\mathbf{n}^{V}\mid\mathbf{n}^{V})).$
The following lemma informs how path-uniqueness simplifies the event $\partial\mathbf{n}^{E}=\partial\mathbf{n}^{V}$.
Lemma 4.5.
Let $\mathbf{n}^{V}\in\mathcal{C}^{V}$ be fixed. Then a current $\mathbf{n}^{E}\in\mathcal{C}^{E}$ satisfies the relation $\partial\mathbf{n}^{E}=\partial\mathbf{n}^{V}$ if and only if $\mathbf{n}^{V}$ is split according to the parities of $\mathbf{n}^{E}$.
Proof.
Suppose first that $\partial\mathbf{n}^{E}=\partial\mathbf{n}^{V}$. Starting from the left, $\mathbf{n}^{E}(i,i+1)$ must have the same parity as $\mathbf{n}^{V}(i)$ because no other edges in $E$ connect to $i$. Proceeding inductively, take $x$ to be such that $\mathbf{n}^{E}(x,x+1)$ has the same parity as $\sum_{y=i}^{x}\mathbf{n}^{V}(y)$. If $\mathbf{n}^{V}(x+1)$ is even, then $\sum_{y=i}^{x+1}\mathbf{n}^{V}(y)$ will have the same parity as $\sum_{y=i}^{x}\mathbf{n}^{V}(y)$, while $\mathbf{n}^{E}(x+1,x+2)$ must have the same parity as $\mathbf{n}^{E}(x,x+1)$ to ensure that an even number of arrivals from $\mathbf{n}^{E}$ touch $x+1$. On the other hand, if $\mathbf{n}^{V}(x+1)$ is odd, then $\sum_{y=i}^{x+1}\mathbf{n}^{V}(y)$ will have the opposite parity as $\sum_{y=i}^{j}\mathbf{n}^{V}(y)$, just as $\mathbf{n}^{E}(x+1,x+2)$ and $\mathbf{n}^{E}(x,x+1)$ must have opposite parities so that an odd number of arrivals from $\mathbf{n}^{E}$ will touch $x+1$. We can thus conclude inductively that $\mathbf{n}^{E}(x,x+1)$ has the same parity as $\sum_{y=i}^{x}\mathbf{n}^{V}(y)$ for all $(x,x+1)$. Arguing symmetrically, the same must be true of $\sum_{y=x+1}^{j}\mathbf{n}^{V}(y)$.
Conversely, given that the parities of $\mathbf{n}^{E}(x,x+1)$, $\sum_{y=i}^{x}\mathbf{n}^{V}(y)$, and $\sum_{y=x+1}^{j}\mathbf{n}^{V}(y)$ are the same for all $(x,x+1)\in E$, fix an arbitrary $x\in V$. For notational convenience, let the expression $\mathbf{n}^{E}(i-1,i)=0$ and $\mathbf{n}^{E}(j,j+1)=0$. It is possible to write
(4.10) $\mathbf{n}^{V}(x)=\sum_{y=i}^{j}\mathbf{n}^{V}(y)-\sum_{y=i}^{x-1}\mathbf{n}^{V}(y)-\sum_{y=x+1}^{j}\mathbf{n}^{V}(y),$
where on the right-hand side, the first summation is even, the second summation has the same parity as $\mathbf{n}^{E}(x-1,x)$, and the third summation has the same parity as $\mathbf{n}^{E}(x,x+1)$. Thus, $\mathbf{n}^{V}(x)$ has the same parity as $\mathbf{n}^{E}(x-1,x)+\mathbf{n}^{E}(x,x+1)$, implying that $x\in\partial\mathbf{n}^{V}$ if and only if $x\in\partial\mathbf{n}^{E}$. It follows that $\partial\mathbf{n}^{E}=\partial\mathbf{n}^{V}$. ∎
This characterization makes it possible to rewrite the denominator as
(4.11) $\begin{split}\mathfrak{P}(\partial\mathbf{n}^{E}=\partial\mathbf{n}^{V}\mid\mathbf{n}^{V})=\prod_{e\in E}&(\mathfrak{P}(\text{$\mathbf{n}^{E}(e)$ even})\mathbbm{1}_{\{\text{$e$ splits $\mathbf{n}^{V}$ evenly}\}}\\ &\qquad+\mathfrak{P}(\text{$\mathbf{n}^{E}(e)$ odd})\mathbbm{1}_{\{\text{$e$ splits $\mathbf{n}^{V}$ oddly}\}}).\end{split}$
We are now ready to prove Lemma 4.3.
Proof of Lemma 4.3.
Simplifying (3.28) with the results of Lemmata 4.3 and 4.5, we obtain
(4.12) $|\operatorname{cov}(\sigma_{i},\sigma_{j})|=\left(\frac{e^{-\sum_{v\in V}h_{v}}\prod_{e\in E}\sqrt{|\mathfrak{P}(\mathbf{n}^{E}(e)\,\text{even})\mathfrak{P}(\mathbf{n}^{E}(e)\,\text{odd})|}}{\prod_{e\in E}(\mathfrak{P}(\text{$\mathbf{n}^{E}(e)$ even})\mathbbm{1}_{\{\text{$e$ splits $\mathbf{n}^{V}$ evenly}\}}+\mathfrak{P}(\text{$\mathbf{n}^{E}(e)$ odd})\mathbbm{1}_{\{\text{$e$ splits $\mathbf{n}^{V}$ oddly}\}})}\right)^{2}.$
For any $s\in E\cup V$ we can observe by series summation that
(4.13) $\mathfrak{P}(\mathbf{n}(s)\,\mathrm{even})=e^{-j_{s}}\cosh(j_{s})$
and
(4.14) $\mathfrak{P}(\mathbf{n}(s)\,\mathrm{odd})=e^{-j_{s}}\sinh(j_{s}),$
so
(4.15) $\frac{\mathfrak{P}(\mathbf{n}^{E}(e)\,\mathrm{even})}{\sqrt{|\mathfrak{P}(\mathbf{n}^{E}(e)\,\mathrm{even})\mathfrak{P}(\mathbf{n}^{E}(e)\,\mathrm{odd})|}}=\sqrt{\coth(|J_{e}|)},$
and
(4.16) $\frac{\mathfrak{P}(\mathbf{n}^{E}(e)\,\mathrm{odd})}{\sqrt{|\mathfrak{P}(\mathbf{n}^{E}(e)\,\mathrm{even})\mathfrak{P}(\mathbf{n}^{E}(e)\,\mathrm{odd})|}}=\operatorname{sign}(J_{e})\sqrt{\tanh(|J_{e}|)},$
and Lemma 4.3 follows. ∎
Our next goal in order to obtain a formula for covariance in terms of the model coefficients will be evaluate the integral over $\mathbf{n}^{V}$. In doing so, we will use an argument similar to the transfer matrix method used to compute the partition function, effectively re-deriving Ising’s transfer matrices in a form more natural when dealing with covariances.
Proof of Theorem 2.4.
Our first goal is to express $\prod_{x=i}^{j-1}f_{(x,x+1)}(\mathbf{n}^{V})$ for fixed $\mathbf{n}^{V}$ in terms of the individual parities of $\mathbf{n}^{V}$ at each vertex. In matrix form, we can write
(4.17) $f_{e}(\mathbf{n}^{V})=\mathtt{v}_{e}^{\intercal}\begin{bmatrix}\sqrt{\coth(|J_{e}|)}&0\\ 0&\operatorname{sign}(J_{e})\sqrt{\tanh(|J_{e}|)}\end{bmatrix}\mathtt{v}_{e},$
where
(4.18) $\mathtt{v}_{e}:=\begin{bmatrix}\mathbbm{1}_{\{\text{$e$ splits $\mathbf{n}^{V}$ evenly}\}}\\ \mathbbm{1}_{\{\text{$e$ splits $\mathbf{n}^{V}$ oddly}\}}\end{bmatrix}.$
Moreover, the fact that any edge splits $\mathbf{n}^{V}$ evenly or oddly if and only if $\sum_{x\in V}\mathbf{n}^{V}(x)$ is even implies that
(4.19) $\mathtt{v}_{e}\mathtt{v}_{e}^{\intercal}=\mathbbm{1}_{\{\text{$\sum_{x\in V}\mathbf{n}^{V}(x)$ is even}\}}\mathrm{I},$
where $\mathrm{I}$ represents the identity matrix. To obtain the recursion, we can also note that
(4.20) $\mathtt{v}_{(x,x+1)}=\begin{bmatrix}\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ even}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ odd}\}}\\ \mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ odd}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ even}\}}\end{bmatrix}\mathtt{v}_{(x-1,x)},$
for any $i<x<j$. To facilitate our argument, we can also add two imaginary edges $(i-1,i)$ and $(j,j+1)$ to the left and right of the lattice, in which case the recursive relation holds for the boundary cases $x\in\{i,j\}$ as well. Using the recursion and multiplying all $\{f_{(x,x+1)}\}_{x=i}^{j-1}$ together yields that
(4.21) $\begin{split}\prod_{x=i}^{j-1}f_{(x,x+1)}(\mathbf{n}^{V})&=\mathtt{v}_{(j,j+1)}^{\intercal}\begin{bmatrix}\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(j)$ even}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(j)$ odd}\}}\\ \mathbbm{1}_{\{\text{$\mathbf{n}^{V}(j)$ odd}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(j)$ even}\}}\end{bmatrix}\\ &\qquad\qquad\cdot\prod_{x=i}^{j}\Bigg{(}\begin{bmatrix}\sqrt{\coth(|J_{e}|)}&0\\ 0&\operatorname{sign}(J_{e})\sqrt{\tanh(|J_{e}|)}\end{bmatrix}\\ &\qquad\qquad\qquad\cdot\begin{bmatrix}\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ even}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ odd}\}}\\ \mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ odd}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ even}\}}\end{bmatrix}\Bigg{)}\mathtt{v}_{(i-1,i)}\\ &\qquad\qquad\cdot\mathbbm{1}_{\{\text{$\sum_{x\in V}\mathbf{n}^{V}(x)$ is even}\}}.\end{split}$
As the events $\sum_{x\in V}\mathbf{n}^{V}(x)$ is even implies that $(i-1,i)$ and $(j,j+1)$ split $\mathbf{n}^{V}$ evenly, the above expression can simplify to
(4.22) $\begin{split}\prod_{x=i}^{j-1}f_{(x,x+1)}(\mathbf{n}^{V})&=e_{1}^{\intercal}\begin{bmatrix}\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(j)$ even}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(j)$ odd}\}}\\ \mathbbm{1}_{\{\text{$\mathbf{n}^{V}(j)$ odd}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(j)$ even}\}}\end{bmatrix}\\ &\qquad\qquad\cdot\prod_{x=i}^{j}\Bigg{(}\begin{bmatrix}\sqrt{\coth(|J_{e}|)}&0\\ 0&\operatorname{sign}(J_{e})\sqrt{\tanh(|J_{e}|)}\end{bmatrix}\\ &\qquad\qquad\qquad\cdot\begin{bmatrix}\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ even}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ odd}\}}\\ \mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ odd}\}}&\mathbbm{1}_{\{\text{$\mathbf{n}^{V}(x)$ even}\}}\end{bmatrix}\Bigg{)}e_{1}.\end{split}$
Then, due to the distributive property and Fubini’s theorem, the integral in the denominator of (4.1) can be decomposed into integrals over the individual values of $\mathbf{n}^{V}(x)$ on vertices $i\leq x\leq j$, yielding Theorem 2.4. ∎
5. Corollaries of the matrix identity
In this section, we aim to show a few results that follow from Theorem 2.4.
Proof of Corollary 2.7.
The denominator of the right-hand side of (2.3) can be bounded by
(5.1) $(e_{1}^{\intercal}Me_{1})^{2}\leq\left\|M\right\|^{2},$
[…]
(5.2) $\left\|M\right\|$ $\leq\prod_{x=0}^{N}\left\|\begin{bmatrix}\sqrt{\coth|J_{x,x+1}|}&0\\ 0&\operatorname{sign}(J_{x,x+1})\sqrt{\tanh|J_{x,x+1}|}\end{bmatrix}\right\|\left\|\begin{bmatrix}\cosh(h_{x})&\sinh(h_{x})\\ \sinh(h_{x})&\cosh(h_{x})\end{bmatrix}\right\|$
(5.3) $=\prod_{x=0}^{N}\sqrt{\coth|J_{x,x+1}|}e^{|h_{x}|}.$
The corollary follows. ∎
Proof of Corollary 2.9.
To prove this result, we recall Lemma 4.3, where the integral in the expression
(5.4) $\mathfrak{E}\left[\prod_{x=i}^{j-1}f_{(x,x+1)}(\mathbf{n}^{V})\right]$
can now be understood properly as an expectation because all coefficients of the lattice are positive. Only $\mathbf{n}^{V}$ such that $\sum_{x=i}^{j}\mathbf{n}^{V}(x)$ is even will make nonzero contributions to the expectation, so it makes sense to condition on this event, so that
(5.5) $\mathfrak{E}\left[\prod_{x=i}^{j-1}f_{(x,x+1)}(\mathbf{n}^{V})\right]=\mathfrak{E}\left[\prod_{x=i}^{j-1}f_{(x,x+1)}(\mathbf{n}^{V})\mid\sum_{x=i}^{j}\mathbf{n}^{V}(x)\,\text{even}\right]\mathfrak{P}\left[\sum_{x=i}^{j}\mathbf{n}^{V}(x)\,\text{even}\right],$
where
(5.6) $\mathfrak{P}\left[\sum_{x=i}^{j}\mathbf{n}^{V}(x)\,\text{even}\right]=e^{-\sum_{x=i}^{j}h_{x}}\cosh\left(\sum_{x=i}^{j}h_{x}\right).$
Due to the result of Lemma 4.5, any $\mathbf{n}^{V}$ with $\sum_{x=i}^{j}\mathbf{n}^{V}(x)$ even specifies a unique set of parities $\boldsymbol{\epsilon}\in\{\mathtt{even},\mathtt{odd}\}^{E}$ such that $\mathbf{n}^{V}$ is split according to $\boldsymbol{\epsilon}$, and it can be seen that the definition of $f_{(x,x+1)}$ for any $x$ only depends on this specified $\mathbf{\epsilon}$. To evaluate the conditional expectation, we can consider these functions to take $\{\mathtt{even},\mathtt{odd}\}^{E}$ as their domain. If we let $\mathtt{odd}\prec\mathtt{even}$ and allow this ordering to extend naturally to a partial ordering on $\{\mathtt{even},\mathtt{odd}\}^{E}$, then the set $\{\mathtt{even},\mathtt{odd}\}^{E}$ form a distributive lattice on which the functions $f_{(x,x+1)}$ are a set of non-negative increasing functions. As a result, the FKG inequality guarantees that
(5.7) $\mathfrak{E}\left[\prod_{x=i}^{j-1}f_{(x,x+1)}(\mathbf{n}^{V})\mid\sum_{x=i}^{j}\mathbf{n}^{V}(x)\,\text{even}\right]\geq\prod_{x=i}^{j-1}\mathfrak{E}\left[f_{(x,x+1)}(\mathbf{n}^{V})\mid\sum_{x=i}^{j}\mathbf{n}^{V}(x)\,\text{even}\right].$
Moreover, as Poisson random variables are always more likely to be even than odd, we have that
(5.8) $\mathfrak{P}((x,x+1)\,\text{splits $\mathbf{n}^{V}$ evenly})\geq\mathfrak{P}((x,x+1)\,\text{splits $\mathbf{n}^{V}$ oddly}),$
implying that
(5.9) $\mathfrak{E}\left[f_{(x,x+1)}(\mathbf{n}^{V})\mid\sum_{x=i}^{j}\mathbf{n}^{V}(x)\,\text{even}\right]\geq\frac{\sqrt{\coth(J_{(x,x+1)})}+\sqrt{\tanh(J_{(x,x+1)})}}{2}.$
With some rearrangement, it can be shown that
(5.10) $\frac{\sqrt{\coth(J_{(x,x+1)})}+\sqrt{\tanh(J_{(x,x+1)})}}{2}=\frac{\sqrt{\coth(J_{(x,x+1)})}}{1+e^{-2J_{(x,x+1)}}}.$
Combining these results with (4.1) yields the upper bound. ∎
Proof of Corollary 2.5.
The statement to be shown can be re-written in the form
(5.11) $\log|\operatorname{cov}_{\{-N,\ldots,N\}}(\sigma_{i},\sigma_{j})|=-2\gamma(\operatorname{dist}(i,j)+o(\operatorname{dist}(i,j))),$
where $o(\operatorname{dist}(i,j))$ represents a term with asymptotic growth that is strictly slower than $\operatorname{dist}(i,j)$ for a nondecreasing sequence $\{N;i,j\}$. It will be more convenient for us to work under the change of basis mentioned in Remark 2.2, so that the left-hand side becomes
(5.12) $\log|\operatorname{cov}_{\{-N,\ldots,N\}}(\sigma_{i},\sigma_{j})|$ $=-2\log\left\{\frac{1}{\sqrt{2}}(e_{1}+e_{2})^{\intercal}\tilde{\mathbf{T}}^{\{i,\ldots,j\}}\frac{1}{\sqrt{2}}(e_{1}+e_{2})\right\}.$
By re-packaging the matrix product on the right, we can re-write
(5.13) $\begin{split}\frac{1}{\sqrt{2}}&(e_{1}+e_{2})^{\intercal}\tilde{\mathbf{T}}^{\{i,\ldots,j\}}\frac{1}{\sqrt{2}}(e_{1}+e_{2})\\ &=\frac{1}{2}\begin{bmatrix}e^{h^{\prime}_{j}}&e^{-h^{\prime}_{j}}\end{bmatrix}\\ &\qquad\cdot\prod_{x=i}^{j-1}\left\{\frac{1}{\sqrt{\sinh(2|J_{(x,x+1)}|)}}\begin{bmatrix}e^{J_{(x,x+1)}}&e^{-J_{(x,x+1)}}\\ e^{-J_{(x,x+1)}}&e^{J_{(x,x+1)}}\end{bmatrix}\begin{bmatrix}e^{h_{x}}&0\\ 0&e^{-h_{x}}\end{bmatrix}\right\}\\ &\qquad\cdot\begin{bmatrix}e^{h^{\prime}_{i}-h_{i}}\\ e^{-h^{\prime}_{i}+h_{i}}\end{bmatrix},\end{split}$
so that if $P$ represents the matrix product from $x=i$ to $j-1$, then we obtain the bounds
(5.14) $\begin{split}\log(\left\|P\right\|_{\ell^{1}})-2-|h^{\prime}_{i}-h_{i}|-|h^{\prime}_{j}|&\leq\log\left\{\frac{1}{\sqrt{2}}(e_{1}+e_{2})^{\intercal}\tilde{\mathbf{T}}^{\{i,\ldots,j\}}\frac{1}{\sqrt{2}}(e_{1}+e_{2})\right\}\\ &\qquad\leq\log(\left\|P\right\|_{\ell^{1}})-2+|h^{\prime}_{i}-h_{i}|+|h^{\prime}_{j}|\end{split}$
The matrices that are multiplied to yield $P$ are all i.i.d. and the integrability conditions $\mathbb{E}[\log(\coth|J_{(0,1)}|)]<\infty$ and $\mathbb{E}|h_{0}|<\infty$ guarantee that the integrability conditions for Oseledets’ multiplicative ergodic theorem are satisfied. The theorem of Oseledets then implies that $\log(\left\|P\right\|_{\ell^{1}})$ will grow, $\nu$-a.s., at a rate given by the upper Lyapunov exponent of the system [Ose65]. Therefore, our remaining task is to prove that the correction of $|h^{\prime}_{i}-h_{i}|+|h^{\prime}_{j}|$ will not alter the exponential rate of decay in almost all cases.
We first may limit our attention to sequences of $\{N;i,j\}$ such that the interval $\{i,\ldots,j\}$ increases from $\{0,1\}$, and consider the set of coefficients for which
(5.15) $|h^{\prime}_{i}-h_{i}|+|h^{\prime}_{j}|\geq\frac{1}{c}\operatorname{dist}(i,j)\qquad\text{infinitely often,}$,
for a given $c\in\mathbb{N}$. Because $\{i,\ldots,j\}$ increases from $\{0,1\}$ it follows that $\operatorname{dist}(i,j)=|i|+|j|$, and based on the observations in Subsection 2.1, it would follow in such cases that
[…]
This can only occur if either $|J_{(i-1,i)}|\geq\frac{|i|}{c}$ or if $|h_{j}|+|J_{(j,j+1)}|\geq\frac{|j|}{c}$ infinitely often. However, because
(5.17) $\sum_{|i|\in\mathbb{N}}\nu\left(|J_{(i-1,i)}|\geq\frac{|i|}{c}\right)$ $=\sum_{|i|\in\mathbb{N}}\nu\left(|J_{(0,1)}|\geq\frac{|i|}{c}\right)$
[…]
(5.19) $=c\mathbb{E}|J_{(0,1)}|$
(5.20) <<$\infty$,< ∞ ,
and similarly
(5.21) $\sum_{|j|\in\mathbb{N}}\nu\left(|h_{j}|+|J_{(j,j+1)}|\geq\frac{|j|}{c}\right)$ $=\sum_{|j|\in\mathbb{N}}\nu\left(|h_{0}|+|J_{(0,1)}|\geq\frac{|j|}{c}\right)$
[…]
(5.23) $=c(\mathbb{E}|h_{0}|+\mathbb{E}|J_{(0,1)}|)$
(5.24) <<$\infty$,< ∞ ,
the Borel–Cantelli Lemma implies that the set of such coefficients is a $\nu$-null set. Taking the union over all $c\in\mathbb{N}_{0}$, we can conclude that the growth of $|h^{\prime}_{i}-h_{i}|+|h^{\prime}_{j}|$ as $\{i,\ldots,j\}$ increases from $\{0,1\}$ must be strictly sublinear for $\nu$-almost all realizations of the coefficients. As the same argument can be made for sequences increasing from any two-integer interval $\{x,x+1\}$, $x\in\mathbb{Z}$, removing a countable union of $\nu$-null sets will ensure that any increasing sequence of $\{N;i,j\}$ will see sublinear growth of $|h^{\prime}_{i}-h_{i}|+|h^{\prime}_{j}|$. We can therefore conclude that $\log|\operatorname{cov}_{\{-N,\ldots,N\}}(\sigma_{i},\sigma_{j})|/\operatorname{dist}(i,j)$ will eventually approach $-2\gamma$ for a full-measure set of parameters. ∎
Acknowledgements
JK acknowledges the financial support of the German Academic Exchange Service (DAAD) as well as the Leibniz Association within the Leibniz Junior Research Group on Probabilistic Methods for Dynamic Communication Networks as part of the Leibniz Competition.
Appendix A Calculations for the truncated chain
Here we provide the necessary calculations to define the truncated chain on the vertices $V^{\prime}:=\{i,\ldots,j\}\subset V$. Recall the notation for the modified Hamiltonian
(A.1) $H^{\prime}(\boldsymbol{\sigma}|V^{\prime})=\sum_{x=i}^{j-1}J_{(x,x+1)}^{\prime}\sigma_{x}\sigma_{x+1}+\sum_{x=i}^{j}h_{x}^{\prime}\sigma_{x},$
for $\boldsymbol{\sigma}|V^{\prime}\in\{\pm 1\}^{V^{\prime}}=:\Omega^{\prime}$ where the coefficients $\{J_{(x,x+1)}^{\prime}\}_{x\in\{i,\ldots,j-1\}}$ and $\{h^{\prime}_{x}\}_{x\in V^{\prime}}$ are chosen such that the measure
(A.2) $\mu^{\prime}(\boldsymbol{\sigma}):=\frac{e^{H^{\prime}(\boldsymbol{\sigma})}}{\sum_{\boldsymbol{\sigma^{\prime}}\in\Omega^{\prime}}e^{H^{\prime}(\boldsymbol{\sigma})}}$
is equal to the restriction of the measure $\mu$ to the $\sigma$-algebra generated by $\{\sigma_{i},\ldots,\sigma_{j}\}$.
Integrating out the values $\boldsymbol{\sigma}|{V\setminus V^{\prime}}:=\{\sigma_{x}\}_{x\in V\setminus V^{\prime}}$, we find
(A.3) $\begin{split}\mu^{\prime}(\boldsymbol{\sigma}|V^{\prime})&=\frac{1}{Z}\sum_{\boldsymbol{\sigma}|{V\setminus V^{\prime}}}e^{h_{-N}\sigma_{-N}+\cdots+J_{(i-1,i)}\sigma_{i-1}\sigma_{i}}\\ &\qquad\qquad\qquad\cdot\,e^{h_{i}\sigma_{i}+\cdots+h_{j}\sigma_{j}}\\ &\qquad\qquad\qquad\cdot\,e^{J_{(j,j+1)}\sigma_{j}\sigma_{j+1}+\cdots+h_{N}\sigma_{N}}\end{split}$
(A.4) $\begin{split}&=\frac{1}{Z}\left(\sum_{\sigma_{-N},\ldots,\sigma_{i-1}}e^{h_{-N}\sigma_{-N}+\cdots+J_{(i-1,i)}\sigma_{i-1}\sigma_{i}}\right)\\ &\qquad\qquad\cdot\,e^{h_{i}\sigma_{i}+\cdots+h_{j}\sigma_{j}}\\ &\qquad\qquad\cdot\,\left(\sum_{\sigma_{j},\ldots,\sigma_{N}}e^{J_{(j,j+1)}\sigma_{j}\sigma_{j+1}+\cdots+h_{N}\sigma_{N}}\right).\end{split}$
As we can always find values $A_{i},B_{i}\in\mathbb{R}$ such that
(A.5) $\log\left(\sum_{\sigma_{-N},\ldots,\sigma_{i-1}}e^{h_{-N}\sigma_{-N}+\cdots+J_{(i-1,i)}\sigma_{i-1}\sigma_{i}}\right)=A_{i}+B_{i}\sigma_{i}$
for $\sigma_{i}=\pm 1$, and we can similarly find $A_{j},B_{j}\in\mathbb{R}$ such that
(A.6) $\log\left(\sum_{\sigma_{j},\ldots,\sigma_{N}}e^{J_{(j,j+1)}\sigma_{j}\sigma_{j+1}+\cdots+h_{N}\sigma_{N}}\right)=A_{j}+B_{j}\sigma_{j}$
for $\sigma_{j}=\pm 1$, it follows that the construction of $H^{\prime}$ is always possible. Moreover, we find that $J^{\prime}_{(x,x+1)}=J_{(x,x+1)}$ for all $i\leq x<j$ and that $h^{\prime}_{x}=h_{x}$ for all $i<x<j$, with the only coefficients that differ from the original model being $h^{\prime}_{i}=h_{i}+B_{i}$ and $h^{\prime}_{j}=h_{j}+B_{j}$. For these values it is apparent that $h^{\prime}_{i}$ and $h^{\prime}_{j}$ depend only on parameters to the left or to the right of $\{i,\ldots,j\}$, respectively, through their dependence on $B_{i}$ and $B_{j}$. In order to bound the correction to $h^{\prime}_{i}$ and $h^{\prime}_{j}$, we also note that
(A.7) $B_{i}$ $=\frac{1}{2}\log\left(\frac{\sum_{\sigma_{-N},\ldots,\sigma_{i-1}}e^{h_{-N}\sigma_{-N}+\cdots+J_{(i-1,i)}\sigma_{i-1}}}{\sum_{\sigma_{-N},\ldots,\sigma_{i-1}}e^{h_{-N}\sigma_{-N}+\cdots-J_{(i-1,i)}\sigma_{i-1}}}\right)$
(A.8) $=\frac{1}{2}\log\left(\frac{\sum_{\sigma_{-N},\ldots,\sigma_{i-1}}e^{h_{-N}\sigma_{-N}+\cdots-J_{(i-1,i)}\sigma_{i-1}}e^{2J_{(i-1,i)}\sigma_{i-1}}}{\sum_{\sigma_{-N},\ldots,\sigma_{i-1}}e^{h_{-N}\sigma_{-N}+\cdots-J_{(i-1,i)}\sigma_{i-1}}}\right)$
(A.9) $\leq\frac{1}{2}\log\left(\frac{e^{2|J_{(i-1,i)}|}\sum_{\sigma_{-N},\ldots,\sigma_{i-1}}e^{h_{-N}\sigma_{-N}+\cdots-J_{(i-1,i)}\sigma_{i-1}}}{\sum_{\sigma_{-N},\ldots,\sigma_{i-1}}e^{h_{-N}\sigma_{-N}+\cdots-J_{(i-1,i)}\sigma_{i-1}}}\right)$
(A.10) $=|J_{(i-1,i)}|,$
and that we can similarly establish that $B_{i}\geq-|J_{(i-1,i)}|$ and that $|B_{j}|\leq|J_{(j,j+1)}|$, yielding the inequalities
(A.11) $|h^{\prime}_{i}-h_{i}|$ $\leq|J_{(i-1,i)}|;$
(A.12) $|h^{\prime}_{j}-h_{j}|$ $\leq|J_{(j,j+1)}|.$
Appendix B Proof of Theorem 2.4 using standard techniques of statistical mechanics
The following proof relies on techniques that are more widely used in statistical mechanics than the random current expansion, including elementary conditional probability and Ising’s transfer matrix method. Like the proof given in Section 4, it requires an observation regarding the structure of the underlying lattice, in this case the fact that the partition function can be decomposed as a product of transfer matrices. Both proofs follow rather naturally once the frameworks and key structural insights have been introduced.
Proof.
We consider the truncated chain for simplicity, using $Z^{\prime}$ to refer to its partition function. As a convenient shorthand, we will define
(B.1) $Z^{\prime}_{++}:=\sum_{\boldsymbol{\sigma}\in\Omega,\sigma_{i}=+1,\sigma_{j}=+1}e^{H^{\prime}(\boldsymbol{\sigma})},$
and similarly define $Z^{\prime}_{+-}$, $Z^{\prime}_{-+}$, and $Z^{\prime}_{--}$. The partition function can be expressed in terms of these quantities as
(B.2) $Z^{\prime}=Z^{\prime}_{++}+Z^{\prime}_{+-}+Z^{\prime}_{-+}+Z^{\prime}_{--}.$
Then, we obtain the expressions
(B.3) $\sigma_{i}\sigma_{j}\right>$ $=\frac{Z^{\prime}_{++}+Z^{\prime}_{--}-Z^{\prime}_{+-}-Z^{\prime}_{-+}}{Z^{\prime}}$
(B.4) $=\frac{(Z^{\prime}_{++}-Z^{\prime}_{+-}-Z^{\prime}_{-+}+Z^{\prime}_{--})(Z^{\prime}_{++}+Z^{\prime}_{+-}+Z^{\prime}_{-+}+Z^{\prime}_{--})}{(Z^{\prime})^{2}}$
and
[…]
Subtracting these expressions will cancel most terms, yielding a simplified way of representing the covariance between the two spins,
(B.6) $\operatorname{cov}(\sigma_{i},\sigma_{j})=\frac{4(Z^{\prime}_{++}Z^{\prime}_{--}-Z^{\prime}_{+-}Z^{\prime}_{-+})}{(Z^{\prime})^{2}}.$
When applying this result to the truncated chain, the expression $Z^{\prime}_{++}Z^{\prime}_{--}-Z^{\prime}_{+-}Z^{\prime}_{-+}$ in the numerator can be identified with the determinant of the transfer matrix product in (1.10). This determinant can be computed explicitly as the product of determinants of the individual transfer matrices. For any $J,h\in\mathbb{R}$, we have
(B.7) $\det\left(\begin{bmatrix}e^{J}&e^{-J}\\ e^{-J}&e^{J}\end{bmatrix}\right)=e^{2J}-e^{-2J}=2\sinh(2J),$
while
(B.8) $\det\left(\begin{bmatrix}e^{h}&0\\ 0&e^{-h}\end{bmatrix}\right)=1,$
implying that
(B.9) $Z^{\prime}_{++}Z^{\prime}_{--}-Z^{\prime}_{+-}Z^{\prime}_{-+}=\prod_{x=i}^{j-1}2\sinh(2J^{\prime}_{x,x+1}).$
The partition function for the truncated chain with free boundary conditions can be expressed as the sum of all entries in the transfer matrix product, which can also be written in the form
(B.10) $Z^{\prime}=\begin{bmatrix}1&1\end{bmatrix}\mathbf{T^{\prime}}_{\{i\cdots j\}}\begin{bmatrix}1\\ 1\end{bmatrix},$
with $\mathbf{T^{\prime}}_{\{i\cdots j\}}$ representing the transfer matrix product obtained by carrying out the construction in (1.10) on the truncated chain. As a result, (B.6) can be repackaged into the form
(B.11) $|\operatorname{cov}(\sigma_{i},\sigma_{j})|=\frac{1}{\left\{\begin{bmatrix}1/\sqrt{2}&1/\sqrt{2}\end{bmatrix}\tilde{\mathbf{T}^{\prime}}_{\{i\cdots j\}}\begin{bmatrix}1/\sqrt{2}\\ 1/\sqrt{2}\end{bmatrix}\right\}^{2}},$
where $\tilde{\mathbf{T}^{\prime}}^{\{i\cdots j\}}$ is the normalized transfer matrix product
(B.12) $\tilde{\mathbf{T}^{\prime}}_{\{i\cdots j\}}:=\begin{bmatrix}e^{h^{\prime}_{j}}&0\\ 0&e^{-h^{\prime}_{j}}\end{bmatrix}\prod_{x=i}^{j-1}\frac{1}{\sqrt{2\sinh(2|J^{\prime}_{x,x+1}|)}}\begin{bmatrix}e^{J^{\prime}_{x,x+1}}&e^{-J^{\prime}_{x,x+1}}\\ e^{-J^{\prime}_{x,x+1}}&e^{J^{\prime}_{x,x+1}}\end{bmatrix}\begin{bmatrix}e^{h^{\prime}_{x}}&0\\ 0&e^{-h^{\prime}_{x}}\end{bmatrix}.$
Using the observation made in Remark 2.2, (B.11) can be transformed into (2.4). ∎